Weston Corporation just paid a dividend of $1.75 a share (i.e., D0 = $1.75). The dividend is expected to grow 11% a year for the next 3 years and then at 5% a year thereafter. What is the expected dividend per share for each of the next 5 years? Do not round intermediate calculations. Round your answers to the nearest cent. D1 = $ D2 = $ D3 = $ D4 = $ D5 = $
Answer #1

D1 = 1.75(1.11) = $1.94

D2 = 1.75(1.11)2 = $2.16

D3 = 1.75(1.11)3 = $2.39

D4 = 1.75(1.11)3(1.05) = $2.51

D5 = 1.75(1.11)3(1.05)2 = $2.64